# CiFlow: Dataflow Analysis and Optimization of Key Switching for Homomorphic Encryption

### v2

@@ IV-D Dataflow Comparison: Arithmetic Intensity
Table II evaluates the off-chip data movement, including $\mathbf{evk}$ and input/output data, of each benchmark. Here, the assumed on-chip memory is 32MB with $\mathbf{evks}$ being streamed on-chip. Across all benchmarks, we see that OC can significantly reduce the total off-chip traffic. The number of operations per HKS benchmark is independent of dataflow. Thus, we can see that the arithmetic intensity (AI) of each is also significantly improved with OC. In a recent study, MAD [25] introduced techniques to optimize and accelerate FHE. Analyzing their optimizations, we find their solution is analogous to our proposed DC dataflow. According to the reported arithmetic intensity in Table II, leveraging the OC dataflow results in 1.43$\times$ to 2.4$\times$ more arithmetic intensity than MP and 1.43$\times$ to 1.98$\times$ more than DC. Furthermore, MAD [25] (and other FHE accelerators) also consider a technique for key compression, which halves the off-chip data movement of keys. Incorporating key compression in our approach will boost our AI to 3.82.
@@ V Methodology
This section provides an overview of our dataflow implementations. We will start by describing the hardware architecture, and then explain the benchmarks used for performance evaluation. Finally, we will introduce the software framework employed for generating HKS instructions for different benchmarks and dataflows.
@@ V-A Hardware Simulation
We implement the dataflows introduced in Section IV for the RPU [14], alongside an extensive design space exploration in Section VI. The RPU is a vector processor supporting Ring-Learning-With-Errors (RLWE)-based algorithms, including HE, with 64, 64-bit vector registers, 64, 64-bit scalar registers, a 32MB vector data memory, and a 1MB scalar data memory, operating at 1.7GHz. These exist alongside an additional modulus register file, which stores the working set of RNS moduli. In our work, the RPU uses 128 lanes. This marks a 2×2$\times$2 × decrease in the size of RNS moduli from the original RPU paper[14], which entailed doubling both the number of vector lanes and vector length to maintain high throughput and keep compute units occupied. We also modified the RPU’s associated instruction set architecture (ISA), $\mathtt{B512}$, to support $1K$ vector length and will refer to it as $\mathtt{B1K}$. $\mathtt{B1K}$ consists of $28$ instructions ranging from general purpose point-wise arithmetic operations to HE-specific shuffle instructions for (i)NTT kernels. Figure 3 presents the RPU microarchitecture, which consists of a front-end to handle instruction fetching, decoding, and control logic, and a backend that contains the high-performance large arithmetic word engines (HPLEs) to efficiently perform our HE-tailored vector operations. To achieve high performance and mask data movement by computation, the RPU includes three distinct queues to fetch independent instructions in parallel. Additionally, the RPU was implemented and verified on a Palladium system [26].
@@ V-B Benchmark Selection
We use the RNS variant of HKS, as presented in Han[17], which optimized the key-switching implementation by leveraging RNS decomposition for enhanced efficiency. The security of HE is defined by $\lambda$, which is a function of $\frac{N}{logQP}$. As mentioned in Cheon[27], 128-bit security is required to defend against dual attacks and to be able to extract the real data from the encrypted output. BTS [15] simulates the multiplication time per slot by sweeping $N$, $L$, and $\lambda$ to identify the optimal configurations for 128-bit security, and refers to these points as BTS1, BTS2, and BTS3. ARK [16] uses a smaller polynomial degree, $2^{16}$, while increasing the number of digits to 4 to satisfy the 128-bit security. Following recent studies [15, 16, 28, 20], we also target 128-bit security and use the latest work parameterization, BTS[15], ARK[16], and DPRIVE [29], to evaluate the performance of our dataflows. Table III summarizes the parameters of the benchmarks used in our research.
@@ V-C Software Framework
To generate HKS code for different benchmarks and dataflows, we break down the HKS algorithm into steps highlighted in Figure 2, and generate instructions for each step independently, based on the $\mathtt{B1K}$ ISA. Our simulation framework includes two distinct tasks, namely memory tasks and compute tasks. Memory tasks handle data transfer between off-chip and on-chip, while compute tasks correspond to steps in the HKS algorithm, executed on the RPU [14]. The challenge of our approach lies in effectively managing the task dependencies. Each task may rely on one or more compute and memory tasks, which must complete their execution before the dependent task can proceed. These dependencies stem from a variety of reasons, such as the need to fetch data from off-chip memory or to create space for subsequent operations using a memory task. Moreover, some tasks depend on data generated by a compute task as part of their input. These dependencies may vary based on the on-chip memory, the specific dataflow, and the benchmarks used. Using the software framework, we generate instructions for each configuration.
@@ V-D Dataflow Simulation
Our goal is to keep as much data as we can on-chip to reduce interactions with off-chip memory. In general, the INTT calculations will be in-place or out-of-place, depending on the on-chip memory. In ModUp_P5, if on-chip space is sufficient, we prioritize storing towers related to $[P_{0}]_{\mathcal{B}}$ and $[P_{1}]_{\mathcal{B}}$, for optimized on-chip memory use and reduced off-chip memory access during subsequent computations. Additionally, ModDown_P2 and ModDown_P3 will begin computations from towers that are already on-chip.
@@ VI Evaluation
In this section, we evaluate HKS performance following the three dataflows described above, using five benchmarks listed in Table III, from recent work, BTS [15], ARK [16], and DPRIVE [29]. The evaluation is done using the RPU [14] as the target hardware platform of our implementations, including the modifications mentioned in Section V-A. While the performance results are specific to the RPU, the general insights and takeaways from the optimizations are broadly applicable. As mentioned earlier, one of the challenges in HKS is properly handling the large amounts of input, intermediate data, and $\mathbf{evks}$. In fact, prior work [15, 16, 30] has found that HE is memory bound. We re-examine this performance limitation considering the different proposed dataflows. First, we assume a large 512MB on-chip SRAM, with 32MB allocated for data and the rest for $\mathbf{evks}$. We sweep bandwidth across all three dataflows to demonstrate the bandwidth saving achievable with OC while maintaining performance equal to MP and DC at higher bandwidths. Next, we eliminate the on-chip SRAM dedicated to storing $\mathbf{evks}$, instead streaming them on-chip, as they are used only once per HKS, and saving substantial on-chip SRAM (up to 16×16$\times$16 ×). We find that a small on-chip memory significantly degrades the performance of a naive implementation (MP), but that our optimizations (OC) result in minimal slowdown. Finally, we increase the computational throughput of the RPU to understand how our proposed dataflow balances the communication-to-compute ratio.
@@ VI-A Dataflow Comparison: Saving Bandwidth
Below we will delve into a detailed comparison between OC, MP, and DC runtimes, taking the MP implementation with 64GB/s off-chip bandwidth and on-chip $\mathbf{evks}$ as our baseline. We will explore the bandwidth at which OC matches baseline performance, denoting it as $OC_{base}$. 64GB/s was chosen as our reference point, given that it represents the peak DDR5 bandwidth in our evaluations, and highlighting cost-effective designs before moving to expensive memory technologies.
@@ VI-A1 DPRIVE & ARK
Figure 4(a) and (d), demonstrate a significant improvement in HKS runtime with OC dataflow. The horizontal line in Figure 4(a) shows that OC operating at 12.8GB/s bandwidth matches the baseline, resulting in a 5$\times$ bandwidth saving. At this specific data point, OC is 2.57$\times$ and 2.96$\times$ faster than DC and MP, respectively. Before becoming compute bound, the RPU is idle for part of its execution time, waiting for dependent memory tasks to be completed. With 12.8GB/s off-chip bandwidth, OC causes the RPU idle for 20.87% of its execution time, outperforming DC and MP with 68.62% and 72.76% idle times, respectively. This highlights OC’s advantages when off-chip bandwidth is a concern and its efficiency in minimizing idle time.
In the case of ARK, the OC implementation with 8GB/s off-chip bandwidth achieves the same performance as MP and DC with 64GB/s bandwidth, resulting in 8$\times$ and 5$\times$ bandwidth savings, respectively. With 8GB/s off-chip bandwidth, OC outperforms MP and DC runtime by notable factors of 4.16$\times$ and 3.22$\times$, and being 2.25$\times$ less idle cycle than MP.
@@ VI-A3 BTS2 & BTS3
By analyzing BTS2 and BTS3, we observe that as the size of the benchmark increases, DC and MP converge due to data expansion in BConv and an increased number of polynomial additions for reduction.
@@
For BTS3, the performance gap between OC and the other two dataflows decreases as the RPU becomes more compute bound, especially compared to BTS2. Despite the decreasing gap between OC and MP/DC, due to the increased input and intermediate data in BTS3, OC still outperforms MP/DC. With 32GB/s bandwidth, OC is 1.3$\times$ faster and 1.7$\times$ less idle than MP and DC, while matching the baseline’s performance, resulting in a 2$\times$ bandwidth saving.
@@ VI-A4 Comparison Across Benchmarks
[…]
Many recent HE algorithmic optimizations, e.g., RNS-based HKS [17], which is used in this paper, require a higher bandwidth. Therefore, focusing on reducing off-chip bandwidth is crucial. The OC dataflow results in 1.30$\times$ to 4.16$\times$ speedup over a simple MP implementation and saves 2$\times$ to 8$\times$ off-chip bandwidth.
@@ VI-B Streaming Evaluation Keys: Trading SRAM for Bandwidth
Analyzing the computations of HKS, the input data and intermediate data (BConv output) are reused for calculating multiple output towers, making it advantageous to keep them on-chip to reduce off-chip communication. On the other hand, the evks are large, ranging from 99MB for DPRIVE to 360MB for BTS3, and only used once per HKS. Consequently, it can be more practical and area efficient to store these large $\mathbf{evks}$ off-chip and load them on-chip in a streaming fashion. To study the effects of streaming $\mathbf{evks}$, we reserve a fraction of off-chip bandwidth and dedicate it to loading the $\mathbf{evks}$ on-chip. In this case, we only have 32MB of on-chip memory to store data and capture on-chip reuse. To determine the required bandwidth allocation for $\mathbf{evks}$, we calculate the ratio of the number of $\mathbf{evks}$ to the number of loaded/stored data.
@@
Figure 5 shows the slowdown for the OC dataflow of each benchmark when streaming $\mathbf{evks}$ from off-chip. The results are shown for two bandwidths per benchmark, as indicated by the clustered bars (e.g., 8 and 23.4 for ARK). The first bandwidth corresponds to the $OC_{base}$ bandwidth specified in Table IV, where the performance is equal to the baseline, assuming $\mathbf{evks}$ to be on-chip. The second bandwidth indicates the required bandwidth to attain equivalent performance when streaming the $\mathbf{evks}$ from off-chip. The slowdown for DPRIVE is less than ARK due to the smaller ratio of $\mathbf{evks}$ to the loaded/stored data (0.66 for ARK and 0.5 for DPRIVE). Among the large benchmarks (BTS1, BTS2, and BTS3), BTS2 has the most slowdown, 1.33$\times$, since that ratio is more than other benchmarks. As shown in Figure 5, with OC dataflow, and 1.7$\times$ (for BTS3) to 2.9$\times$ (for ARK) more bandwidth we can achieve the same performance as storing the $\mathbf{evks}$ on-chip while saving 16$\times$ on-chip memory. However, by storing $\mathbf{evks}$ off-chip and using OC dataflow with 54.6GB/s bandwidth, BTS3 still achieves the same performance as the baseline (MP with 64GB/s and $\mathbf{evks}$ on-chip), and saves 1.1$\times$ bandwidth and 16$\times$ on-chip memory. In conclusion, storing $\mathbf{evks}$ off-chip saves 16$\times$ on-chip memory and by implementing OC dataflow, we can save up to 3.3$\times$ bandwidth to have the same performance as the baseline. Additionally, storing $\mathbf{evks}$ off-chip and keeping 32MB on-chip memory for data, decreases the RPU area from $521.85mm^{2}$ to $41.85mm^{2}$.

### v4

@@ IV-D Dataflow Comparison: Arithmetic Intensity
Table II evaluates the off-chip data movement, including $\mathbf{evks}$ and input/output data, of each benchmark. Here, the assumed on-chip memory is 32MB with $\mathbf{evks}$ being streamed on-chip. Across all benchmarks, we see that OC can significantly reduce the total off-chip traffic. The number of operations per HKS benchmark is independent of dataflow. Thus, we can see that the arithmetic intensity (AI) of each is also significantly improved with OC.
In a recent study, MAD [26] introduced techniques to optimize and accelerate FHE. Analyzing their optimizations, we find their solution is analogous to our proposed DC dataflow. According to the reported arithmetic intensity in Table II, leveraging the OC dataflow results in 1.43$\times$ to 2.4$\times$ more arithmetic intensity than MP and 1.43$\times$ to 1.98$\times$ more than DC. Furthermore, MAD [26] (and other FHE accelerators) also consider a technique for key compression, which halves the off-chip data movement of keys. Incorporating key compression in our approach will further boost our AI to 3.82.
@@ V Methodology
This section provides an overview of our dataflow implementations. We will start by describing the hardware architecture and then explain the benchmarks used for performance evaluation. Finally, we will introduce the software framework employed for generating HKS instructions for different benchmarks and dataflows.
@@ V-A Hardware Simulation
We implement the dataflows introduced in Section IV on the RPU [15] and conduct an extensive evaluation in Section VI. The RPU is a vector processor supporting Ring-Learning-With-Errors (RLWE)-based algorithms, including HE, with 64, 64-bit vector registers, 64, 64-bit scalar registers, a 32MB vector data memory, and a 1MB scalar data memory, operating at 1.7GHz. These exist alongside an additional modulus register file, which stores a set of RNS moduli. Figure 3 presents the RPU microarchitecture. It consists of a front-end to handle instruction fetching, decoding, and control logic, and a backend that contains the high-performance large arithmetic word engines (HPLEs) to efficiently perform HE operations. A parameterized simulator is used to test different RPU configurations (e.g., SRAM capacity and computational throughput), and supports arbitrarily large parameter sets.
This work uses a 2×2$\times$2 × smaller RNS moduli compared to the original RPU paper [15]. Here, we assume 128 HPLEs and have modified the RPU’s associated instruction set architecture (ISA), $\mathtt{B512}$, to support a vector length of $1K$, referred to as $\mathtt{B1K}$, to maintain high throughput and keep compute units occupied. Longer vectors make hardware efficient, e.g., taking pressure off the frontend and improving compute utilization. $\mathtt{B1K}$ consists of $28$ instructions ranging from general purpose point-wise arithmetic operations to HE-specific shuffle instructions for (i)NTT kernels.
Since FHE is data-oblivious, all memory addresses are known at compile time, and the behavior is independent of input values. This property enables data prefetching through decoupling, where the compiler can order memory requests to overlap data movement with computation. This optimization leverages the deeply decoupled microarchitecture of the RPU, which employs three distinct queues to fetch independent compute, shuffle, and memory instructions in parallel. This overlap helps mask latency and improve overall performance.
@@ V-B Benchmark Selection
We use the RNS variant of HKS, as presented in Han [18], which optimizes the key-switching implementation by leveraging RNS decomposition for enhanced efficiency. The security of HE is defined by $\lambda$, which is a function of $\frac{N}{logQP}$. As mentioned in Cheon [27], 128-bit security is required to defend against dual attacks and to be able to extract the real data from the encrypted output. BTS [16] simulates the multiplication time per slot by sweeping $N$, $L$, and $\lambda$ to identify the optimal configurations for 128-bit security, and refers to these points as BTS1, BTS2, and BTS3. ARK [17] uses a smaller polynomial degree, $2^{16}$, while increasing the number of digits to 4 to satisfy the 128-bit security. As in recent studies [16, 17, 28, 21] we provide 128-bit security using proper FHE parameterization from BTS[16], ARK[17], and DPRIVE [29], to evaluate the performance of our dataflows. Table III summarizes the parameters of the benchmarks used in our research.
@@ V-C Software Framework
To generate HKS code for different benchmarks and dataflows, we break down the HKS algorithm into steps highlighted in Figure 2 and generate instructions for each step independently, based on the $\mathtt{B1K}$ ISA. Our simulation framework includes two distinct tasks: memory and compute. Memory tasks handle data transfer between off-chip and on-chip, while compute tasks correspond to steps in the HKS algorithm, executed on the RPU [15]. The challenge of our approach lies in effectively managing task dependencies and the sequence of instructions. Each task may rely on one or more compute and memory tasks, which must complete their execution before the dependent task can proceed. These dependencies stem from a variety of reasons, such as the need to fetch data from off-chip memory or to create space for subsequent operations using a memory task. Moreover, some tasks depend on data generated by a compute task as part of their input. These dependencies may vary based on the on-chip memory, the specific dataflow, and the benchmarks used. Using the software framework, we generate instructions for each configuration and dataflow and define their dependencies. The framework has two distinct queues, one for memory tasks and one for compute tasks. The tasks at the front of each queue are fetched and executed in parallel once all the task’s dependencies are resolved. If there are no dependencies between a memory task and a compute task, the off-chip data movement can be masked by the computation on the RPU.
@@ V-D Dataflow Simulation
Our goal is to reduce interactions with off-chip memory by increasing on-chip data reuse. In general, when sufficient on-chip memory is available, INTT outputs can be stored on-chip and reused. However, with limited on-chip storage, the INTT output must be stored off-chip and reloaded for subsequent computations, increasing the off-chip data movement. For example, in BTS3, using the MP implementation, INTT is applied to all 45 input towers, leaving no space for storing all INTT outputs on-chip. In contrast, using OC, INTT is applied to 30 towers, allowing the INTT outputs to be stored on-chip and used for later computation.
In ModUp_P5, if on-chip space is sufficient, we prioritize storing towers related to $[P_{0}]_{\mathcal{B}}$ and $[P_{1}]_{\mathcal{B}}$, for optimized on-chip memory use and reduced off-chip memory access during subsequent computations. Additionally, ModDown_P2 and ModDown_P3 will begin computations from towers that are already on-chip.
@@ VI Evaluation
In this section, we evaluate HKS performance following the three dataflows described above, using five benchmarks listed in Table III from recent work (BTS [16], ARK [17], and DPRIVE [29]). The evaluation is done using the RPU [15] as the target hardware platform of our implementations, including the modifications mentioned in Section V-A. While the performance results are specific to the RPU, the general insights and takeaways from the optimizations are broadly applicable. As mentioned earlier, one of the challenges in HKS is properly handling large amounts of input, intermediate data, and $\mathbf{evks}$. In fact, prior work [16, 17, 30] has found that HE is memory bound. We re-examine this performance limitation considering the different proposed dataflows. First, we assume a large 392MB on-chip SRAM, with 32MB allocated for data and the rest for $\mathbf{evks}$, pre-loaded on-chip. We sweep bandwidth across all three dataflows to demonstrate the bandwidth saving achievable with OC while maintaining performance equal to MP and DC at higher bandwidths. Next, we eliminate the on-chip SRAM dedicated to storing $\mathbf{evks}$, instead streaming them on-chip, as they are used only once per HKS, and saving substantial on-chip SRAM (up to 12.25×12.25$\times$12.25 ×). We find that a small on-chip memory significantly degrades the performance of a naive implementation (MP), but that our optimizations (OC) result in minimal slowdown. Finally, we increase the computational throughput of the RPU to understand how our proposed dataflow balances the communication-to-compute ratio.
@@ VI-A Dataflow Comparison: Saving Bandwidth
Below we will delve into a detailed comparison between OC, MP, and DC runtimes, taking the MP implementation with 64GB/s off-chip bandwidth and on-chip pre-loaded $\mathbf{evks}$ as our baseline. We will explore the bandwidth at which OC matches baseline performance, denoting it as $OC_{base}$. 64GB/s was chosen as our reference point, given that it represents the peak DDR5 bandwidth in our evaluations, and highlighting cost-effective designs before moving to expensive memory technologies.
@@ VI-A1 DPRIVE & ARK
Figure 4(a) and (d), demonstrate a significant improvement in HKS runtime with OC dataflow. The horizontal line in Figure 4(a) shows that OC operating at 12.8GB/s bandwidth matches the baseline, resulting in a 5$\times$ bandwidth saving. At this specific data point, OC is 2.57$\times$ and 2.96$\times$ faster than DC and MP, respectively. Before becoming compute bound, the RPU is idle for part of its execution time, waiting for dependent memory tasks to be completed. With 12.8GB/s off-chip bandwidth, OC causes the RPU to be idle for 20.87% of its execution time, outperforming DC and MP with 68.62% and 72.76% idle times, respectively. This highlights OC’s advantages when off-chip bandwidth is a concern and its efficiency in minimizing idle time.
In the case of ARK, the OC implementation with 8GB/s off-chip bandwidth achieves the same performance as MP and DC with 64GB/s bandwidth, resulting in 8$\times$ and 5$\times$ bandwidth savings, respectively. With 8GB/s off-chip bandwidth, OC outperforms MP and DC runtime by notable factors of 4.16$\times$ and 3.22$\times$ and being 2.25$\times$ less idle cycle than MP.
@@ VI-A3 BTS2 & BTS3
By analyzing BTS2 and BTS3, we observe that as the benchmark size increases, DC and MP converge due to the large data size during BConv expansion and more polynomial additions for reduction.
@@
For BTS3, the performance gap between OC and the other two dataflows decreases as the RPU becomes more compute bound. Despite the decreasing gap between OC and MP/DC, due to the increased input and intermediate data in BTS3, OC still outperforms MP/DC. With 32GB/s bandwidth, OC is 1.3$\times$ faster and 1.7$\times$ less idle than MP and DC, while matching the baseline’s performance, resulting in a 2$\times$ bandwidth saving.
@@ VI-A4 Comparison Across Benchmarks
Table IV summarizes the previous discussion, with “$OC_{base}$” indicating the bandwidth where OC matches the baseline performance (MP with 64MB/s bandwidth), and “Saved BW” showing the bandwidth savings achieved by OC. Runtimes are reported based on the $OC_{base}$ bandwidth. Notably, BTS1 has less speedup due to its larger size compared to ARK and DPRIVE and more data movement for the reduction than BTS2 and BTS3. This can also be seen in Table II, where the AI (Arithmetic Intensity) improvement for BTS1 in OC is less than other benchmarks.
Many recent HE algorithmic optimizations, e.g., RNS-based HKS [18], which is used in this paper, require a higher bandwidth. Therefore, reducing off-chip bandwidth is crucial. The OC dataflow results in 1.30$\times$ to 4.16$\times$ speedup over a simple MP implementation and saves 2$\times$ to 8$\times$ off-chip bandwidth.
@@ VI-B Streaming Evaluation Keys: Trading SRAM for Bandwidth
Analyzing the computations of HKS, the input data and intermediate data (BConv output) are reused for calculating multiple output towers, making it advantageous to keep them on-chip to reduce off-chip communication. On the other hand, the evks are large, ranging from 99MB (DPRIVE) to 360MB (BTS3), and are only used once per HKS. Consequently, it can be more practical and area efficient to store these large $\mathbf{evks}$ off-chip and load them on-chip in a streaming fashion. To study the effects of streaming $\mathbf{evks}$, we reserve a fraction of off-chip bandwidth and dedicate it to loading the $\mathbf{evks}$ on-chip. In this case, we only have 32MB of on-chip memory to store data and capture on-chip reuse. To determine the required bandwidth allocation for $\mathbf{evks}$, we calculate the ratio of the number of $\mathbf{evks}$ to the number of loaded/stored data.
Figures 5 and  6 illustrate the effects of storing $\mathbf{evks}$ off-chip and HKS runtime as a function of bandwidth. The dotted lines represent the runtime with pre-loaded $\mathbf{evks}$ stored on-chip. As shown in the figures, storing $\mathbf{evks}$ off-chip maintains the same trend but with shifted absolute values due to the increased bandwidth pressure from streaming $\mathbf{evks}$. According to Table IV, for BTS3, 32GB/s is the $OC_{base}$ bandwidth (equal performance as baseline). However, Figure 5 demonstrates that by streaming $\mathbf{evks}$ on-chip, the OC dataflow can achieve baseline performance with a higher bandwidth of 45.62GB/s. Similarly, Figure 6 shows this concept for ARK, where a bandwidth of 23.4GB/s allows ARK to match baseline performance. While in both these cases more bandwidth is required to match performance, we argue the increase is minor compared to the 12.25×12.25$\times$12.25 × reduction in on-chip SRAM.
Figure. 5: HKS runtime for BTS3 with $\mathbf{evks}$ being off-chip.
Figure. 6: HKS runtime for ARK with $\mathbf{evks}$ being off-chip.
@@
Figure 7 shows the slowdown for the OC dataflow of each benchmark when streaming $\mathbf{evks}$ from off-chip. The results are shown for two bandwidths per benchmark, as indicated by the clustered bars (e.g., 8 and 23.4 for ARK). The first bandwidth corresponds to the $OC_{base}$ bandwidth specified in Table IV, where the performance is equal to the baseline, assuming $\mathbf{evks}$ to be on-chip. The second bandwidth indicates the required bandwidth to attain equivalent performance when streaming the $\mathbf{evks}$ from off-chip. The slowdown for DPRIVE is less than ARK due to the smaller ratio of $\mathbf{evks}$ to the loaded/stored data (0.66 for ARK and 0.5 for DPRIVE). Among the large benchmarks (BTS1, BTS2, and BTS3), BTS2 has the most slowdown, 1.33$\times$, since the mentioned ratio is more than other benchmarks. As shown in Figure 7, with the OC dataflow, and 1.3$\times$ (BTS1) to 2.9$\times$ (ARK) more bandwidth we can achieve the same performance as storing the $\mathbf{evks}$ on-chip while saving 12.25$\times$ on-chip memory. However, compared to the original 64GB/s MP implementation with $\mathbf{evks}$ on-chip, the OC dataflow with keys off-chip still saves 1.4$\times$ up to 3.3$\times$ bandwidth for BTS3 and BTS2, respectively, to achieve the same performance. In conclusion, storing $\mathbf{evks}$ off-chip saves 12.25$\times$ on-chip memory and by implementing OC dataflow, we can save up to 3.3$\times$ bandwidth to have the same performance as the baseline. Additionally, storing $\mathbf{evks}$ off-chip and keeping 32MB on-chip memory for data, decreases the RPU area from $401.85mm^{2}$ to $41.85mm^{2}$.